OpenStudy (anonymous):

Sid is a cowboy. He lives on the coordinate point (10,5). He has been riding his horse Abbey all day. He is tired and wants to get a drink of water. There is a river flowing across the x-axis. His horse is also tired so after drinking some water Sid wants to go to the horse stall located at (2,10). What is the shortest distance that Sid has to cover to complete all his tasks?

OpenStudy (jdoe0001):

 \(\large {\text{distance between 2 points}\\ d = \sqrt{(x_2-x_1)^2 + (y_2-y_1)^2}}\)

OpenStudy (anonymous):

ok thanks

OpenStudy (jdoe0001):

\(\begin{array}{llll} x_1&y_1&x_2&y_2\\ (10,&5)\qquad (&2,&10) \end{array} \\ \quad \\\bf d = \sqrt{(x_2-x_1)^2 + (y_2-y_1)^2}\\ \quad \\ \implies d = \sqrt{((2)-(10))^2 + ((10)-(5))^2}\) yw